# Deciphering Textual Authenticity: A Generalized Strategy through the Lens of Large Language Semantics for Detecting Human vs. Machine-Generated Text

### v1

@@ 7.7 Adversarial Robustness Experiment
AI Writer Article Forge Kafkai Reddit Bot
T5LLMCipher-MC -6.1 -15.4 -13.7 -16.4
T5LLMCipher-Bi -3.1 -21.4 -17.0 -20.1
T5LLMCipher-C-KNN -9.6 -6.3 -16.5 -25.3
T5LLMCipher-KNN -7.6 -6.7 -19.7 -18.5
Table 8: Experiment comparing the percentage change in recall from unperturbed samples to adversarially perturbed samples. Lower values indicate better adversarial robustness.
@@
To demonstrate our method’s robustness to adversarial attacks, we evaluate our system against a recently introduced perturbation attack method for evading machine-generated text detection, DFTFooler [41]. DFTFooler carefully modifies machine-generated text to evade detection while attempting to preserve linguistic quality. This is achieved by pinpointing words in the text which are confidently predicted by a pre-trained LLM and substituting them with synonyms that are semantically similar yet less confidently predicted. The synonym selection process adheres to a meticulous protocol to uphold linguistic quality: First, potential synonyms are extracted based on cosine similarity between word embeddings, ensuring a semantic alignment with the original word. Subsequent steps involve part-of-speech checking, to maintain grammatical congruity, and semantic similarity checking to ensure the substituted word does not deviate semantically from the original text. The selected synonym is then one that is predicted with lower confidence by the language model, which alters the predictability of the word choices in the text. Notably, unlike other adversarial techniques, DFTFooler achieves this without necessitating any queries to the detection model, which makes it an attractive choice for potential adversaries who would want to attack a black-box detection model [23]. We use machine-generated texts perturbed texts using DFTFooler, setting the low probability threshold to 0.01, the synonym similarity threshold to 0.7, and allowing for 10 word perturbations. We show the results of this attack on the same datasets of in-the-wild data as in Section 7.6 on the T5LLMCipher-MC, T5LLMCipher-Bi, T5LLMCipher-C-KNN and T5LLMCipher-KNN classifiers. We randomly sample 200 texts from the human-generated texts and 200 texts from the machine-generated texts from each of the four datasets for a total of 1600 samples. We show the results of this experiment in Table 8. Similar to the experiment in Table 2, we present the percent reduction in recall, where lower numbers indicate greater adversarial robustness. These results show that the T5LLMCipher-MC classifier was the most robust model, with an average percent reduction in recall of just 12.9%. Compared to the results in Table 2, where the top performing model on unseen data, RoBERTa-D, showed an average percent reduction in recall of 73.8% on adversarially perturbed unseen data, our approach shows a superior robustness to adversarial attacks.
@@ 8 Discussion and Future Work
While the T5LLMCipher-MC classifier showed the best adversarial robustness in Table 8, it’s worth emphasizing that this heightened robustness wasn’t exclusive to this model. Across the board, models utilizing the embeddings from the LLM encoder displayed a consistent pattern of formidable resistance to adversarial attacks. This underscores the intrinsic robustness of the LLM embeddings themselves. Such robustness implies that the embeddings capture fundamental and semantically rich representations of the text, making it challenging for adversarial perturbations to meaningfully distort the text in a way that would deceive our classifiers. This inherent strength of LLM embeddings augments the credibility and reliability of our detection models, especially in real-world scenarios where adversarial attacks are becoming increasingly sophisticated.
The results of the text generator attribution experiment in Table 4 and Table 3 show that it is possible to not only distinguish between human and machine-generated text, but it is also possible to attribute the text to a specific generator. Identifying the source generator of machine-generated text offers an array of potential applications and benefits in our increasingly digital-centric world. Firstly, in media and journalism, such identification can serve as a mechanism to verify the authenticity of content. Just as systems like the recently introduced LAMBRETTA are designed to flag false information on social media platforms for improved content moderation  [39], identifying the origin of machine-produced texts can help curb the spread of misinformation and ensure the credibility of sources. From a platform regulation and moderation standpoint, tracking down the origin of automated content can assist in managing the flood of posts, flagging suspicious activity, or prioritizing human interactions  [14]. In educational settings, the ability to discern between student-generated content and AI-produced text might prove crucial in maintaining academic integrity. Additionally, for researchers and developers, understanding the nuances and biases of different generators can aid in refining models, and benchmarking performance.
@@
In this study, our primary focus was to emphasize the efficacy of LLM encoder embeddings in detecting machine-generated text, rather than on the complexity of the classifiers. Notably, our results show state-of-the-art generalization and robustness against adversarial attacks, despite the relatively simple nature of the MLP and KNN classifiers. This underscores the inherent quality and potential of the embeddings themselves towards the task of detecting machine-generated text. Given these promising outcomes with simple models, future research that explores more sophisticated architectures in tandem with the LLM encoder embeddings holds significant potential to push generalizability and robustness even further.

### v3

@@ 7.7 Adversarial Robustness Experiment
AI Writer Article Forge Kafkai Reddit Bot Avg
T5LLMCipher-MC 54.9 77.3 54.9 81.5 67.2
T5LLMCipher-Bi 62.4 66.9 22.8 50.0 50.5
T5LLMCipher-C-KNN 31.0 53.0 22.4 53.1 39.9
T5LLMCipher-KNN 19.8 23.7 37.3 37.3 29.5
BERT-D 54.6 55.2 52.1 58.8 55.2
GLTR-BERT 29.7 65.9 10.9 73.0 44.9
GLTR-GPT2 9.1 57.3 10.8 76.9 38.5
RoBERTa-D 13.0 36.6 29.1 82.6 40.3
RADAR 49.6 64.3 58.9 36.2 52.3
Fast-DetectGPT 29.1 54.0 54.0 97.7 58.7
Table 8: Experiment comparing F1 score with respect to machine-generated text of existing state-of-the art approaches to our method on adversarially perturbed in-the-wild text.
@@
To demonstrate our method’s robustness to adversarial attacks, we evaluate our system against a recently introduced perturbation attack method for evading machine-generated text detection, DFTFooler [47]. DFTFooler carefully modifies machine-generated text to evade detection while attempting to preserve linguistic quality. This is achieved by pinpointing words in the text which are confidently predicted by a pre-trained LLM and substituting them with synonyms that are semantically similar yet less confidently predicted. The synonym selection process adheres to a meticulous protocol to uphold linguistic quality: First, potential synonyms are extracted based on cosine similarity between word embeddings, ensuring a semantic alignment with the original word. Subsequent steps involve part-of-speech checking, to maintain grammatical congruity, and semantic similarity checking to ensure the substituted word does not deviate semantically from the original text. The selected synonym is then one that is predicted with lower confidence by the language model, which alters the predictability of the word choices in the text. Notably, unlike other adversarial techniques, DFTFooler achieves this without necessitating any queries to the detection model, which makes it an attractive choice for potential adversaries who would want to attack a black-box detection model [26]. We use machine-generated perturbed texts using DFTFooler, setting the low probability threshold to 0.01, the synonym similarity threshold to 0.7, and allowing for 10 word perturbations. We show the results of this attack on the same datasets of in-the-wild data as in Section 7.6 on the T5LLMCipher-MC, T5LLMCipher-Bi, T5LLMCipher-C-KNN and T5LLMCipher-KNN classifiers, as well as test this attack on the other state-of-the art classifiers tested in Section 7.6. We randomly sample 200 texts from the human-generated texts and 200 texts from the machine-generated texts from each of the four datasets for a total of 1600 samples. We show the results of this experiment in Table 8, where we report the F1 score with respect to the machine-generated text class. These results show that the T5LLMCipher-MC classifier was the most adversarially robust model, with an F1 score of 67.2. This represents a reduction of just 6.1 points from the non-adversarially perturbed in-the-wild text from the experiment in Table 7.
@@ 8 Discussion and Future Work
The results of the text generator attribution experiment in Table 4 and Table 3 show that it is possible to not only distinguish between human and machine-generated text, but it is also possible to attribute the text to a specific generator. Our framework successfully attributes authorship to specific generators included in the T5LLMCipher training data. However, for content produced by machine text generators not included in the initial training, while our system can recognize the content as not human-generated, it cannot pinpoint the exact generator. Detecting a novel generator necessitates retraining of T5LLMCipher for accurate attribution. Fortunately, this process is computationally lightweight.
@@
In this study, our primary focus was to emphasize the efficacy of LLM encoder embeddings in detecting machine-generated text, rather than on the complexity of the classifiers. Notably, our results show state-of-the-art generalization and robustness against adversarial attacks, despite the relatively simple nature of the MLP and KNN classifiers. This underscores the inherent quality and potential of the embeddings themselves towards the task of detecting machine-generated text. Given these promising outcomes with simple models, future research that explores more sophisticated architectures in tandem with the LLM encoder embeddings holds significant potential to push generalizability and robustness even further. Additionally, adversarial training strategies could potentially be utilized to improve adversarial robustness of classifiers built on top of these embeddings.
8.1 Ethical Considerations
The deployment of human versus machine text detection systems warrants important ethical considerations when deployed, particularly in light of the potential for false positives where genuine human-written texts are mistakenly identified as machine-generated. Such misclassifications carry significant ramifications, including but not limited to, the erosion of trust in automated systems, potential harm to individuals’ reputations, and the broader societal impact of inadvertently censoring legitimate expressions in digital communications. In educational contexts, misattributed authorship due to AI-generated content can lead to significant consequences, as seen in an incident where a professor threatened to fail his entire class based on a machine-text detection tool [60]. This underscores the critical need for caution in relying on automated detectors for decision-making. Such tools are not foolproof and may lead to unjust outcomes, making it imperative to use them with caution. Machine-text detection tools have also been shown to be biased against non-native English writers [37], adding even more difficulty in achieving accurate and fair evaluations of texts. In this context, it is crucial to emphasize the limitations of relying solely on systems that assess textual authenticity based on input text alone. Therefore, it is important to recognize that automated machine-text detection systems should not be used for autonomous decision making without human supervision, especially in high-stakes scenarios.
@@ 9 Conclusion
Acknowledgments
This research project and the preparation of this publication were funded in part by the Department of Homeland Security (DHS), United States Secret Service, National Computer Forensics Institute (NCFI) via contract number 70US0920D70090004 and by NSF Grant No. 2245983.
Appendix A Dataset Details
Domain Human Semantic Similarity $\uparrow$ Human Stylistic Similarity $\uparrow$ USE-Cosine Similarity $\uparrow$
Arxiv 5.52 5.66 0.67
Peerread 3.88 4.54 0.52
Reddit 4.53 4.28 0.65
Wikihow 5.70 4.94 0.68
Wikipedia 4.44 4.36 0.57
Table 9: The average semantic similarity and stylistic similarity score given by the human evaluations across 5 domains. "USE-Cosine Similarity" is the average cosine similarity of the pairs of human and machine texts after using the Universal Sentence Encoder (USE) to create vector representation of the texts, where a value of 0 means no semantic similarity and 1 means perfect semantic similarity.
A.1 Prompts
The prompts given to the LLMs to create machine-generated text varied by domain. The PeerRead Reviews domain used four different prompts, to generate a set of peer reviews with different perspectives similar to a set of human generated peer reviews. Each prompt was provided to the LLM independently of the other prompts, and each prompt contains information from which to generate text similar to the human written text that the machine text is paired with. The information unique to each individual prompt is illustrated in [brackets] for illustrative purposes. Below are examples of prompts used for each domain.
  • Wikipedia: Prompts included the title of the Wikipedia article. Example Prompt: Write a Wikipedia article with the title [title], the article should at least have 250 words.
  • Reddit ELI5: Prompts included the question asked in the Reddit post. Example Prompt: Write a couple of paragraphs of at least 500 characters each to address the following question. [question].
  • WikiHow: Prompts included the title and headline in the Wikihow article. Example Prompt: Write a WikiHow article content given a title and a headline. Use approximately 300 words. Title: [title] Headline: [headline].
  • PeerRead Reviews: Prompts included the title and abstract of the research paper. Example Prompt: Write a peer review by first describing what problem or question this paper addresses, then strengths and weaknesses, for the paper [title], its main content is as below: [abstract].
  • Arxiv Abstract: Write a long abstract of a scientific paper from arXiv.org. Use approximately 200-400 words. Title: [title]. Abstract:
A.2 Dataset Assumptions
Although it is not possible to assure with absolute certainty that each entry within the human-annotated datasets was solely produced by human, it is important to note that all data were compiled before the widespread availability of sophisticated and easily accessible language models capable of producing human-like text. The most recent collection of human-written texts in the dataset we use date back to early 2022, predating the introduction of ChatGPT or Gemini.
To ensure that the machine-generated text was reasonably similar to the human-generated texts that the machine-generated texts were based on, we use the Universal Sentence Encoder (USE)  [20] to measure cosine similarity, and human evaluations of style and semantic similarity. USE is first utilized to create high-dimensional vector representations of the texts. Then we measured the cosine similarity between the human-machine text pairs. This method is used to measure the semantic similarity of the human and machine text as an automatic evaluator. Cosine similarity values of 1 indicate a perfect semantic similarity, and values of 0 indicate no semantic similarity. In the human evaluations of text similarity, a 7-point Likert Scale test is used to measure the participant’s agreement or disagreement on a scale from 1 (strongly disagree) to 7 (strongly agree) with the statements of "The two text samples are similar in style?" and "The two text samples are similar in semantics". The authors were used to evaluate the similarity of 250 human-machine text-pairs, and their scores on these statements across these text-pairs were averaged. The 250 text-pairs were randomly sampled to include 10 text-pairs across each of the combinations of five domains and five generators in the M4 dataset described in Section 4. We report the average human evaluation scores across domains in Table 9. We also report the average cosine similarity across domains on the same samples used for human evaluations in Table 9. The average USE cosine similarity across all the evaluated samples is 0.62, which indicates that the level of semantic similarity is moderately high based on this metric. The median sentence length of the human generated and machine generated texts is 295 words and is 252 words, respectively. Given the lengths of the text samples and the expected decrease in similarity measures for longer texts generated by LLMs, the observed average similarity of 0.62 indicates a significant resemblance between the text pairs. This finding is further supported by human evaluation, as presented in Table 9.
As seen in Table 9, the average scores on stylistic and semantic similarity of the human and machine texts indicate good quality of the generated texts, with the Arxiv domain having the highest average human evaluation scores of 5.52 and 5.66 on semantic similarity and stylistic similarity, respectively. The level of similarity of texts between human and machine text pairs varied by domain, but were fairly consistent nonetheless. For instance, in PeerRead Reviews, there were bigger discrepancies between the human-generated and machine-generated samples than in other domains. One potential reason for this mismatch in semantics in the PeerRead Reviews dataset is that the machine generated peer reviews do not have access to the main body of the research papers and only consider the title and abstract of the paper, whereas the human reviewers have access to the full body of the paper in addition to the title and abstract. However, in general, the quality of the generated text seemed to be on par with the human text, even though the actual semantics did not always match on some of the datasets. We acknowledge that the evaluation of the human and machine texts in this experiment could be biased due to the author-led evaluations that might favor interpretations supporting the similarity of the texts. However, this limitation’s impact on our overall system is likely minimal due to several factors. First, the USE cosine similarity measurement corroborates the author-led evaluations, providing automated confirmation of the similarity of the samples. Additionally, the model’s strong performance on in-the-wild data as shown in Table 7 shows that the impact of this limitation on the validity of our model’s results is minimal.